Defn: f is a function if Examples: Prove that s(x) = ±\sqrt{x} is not a function: Defn: f is an even function if Examples: Prove that f(x) = x is not an even function: Defn: f is an odd function if Examples: Prove that f(x) = 1 is not an odd function:

Defn: f is an increasing function on the interval [a, b] if Examples: Prove that f(x) = x^{2} is not an increasing function on the interval [-1, 0]: Defn: f is an decreasing function on the interval [a, b] if Examples: Prove that f(x) = x^{2} is not an decreasing function on the interval [0, 1]: